For a certain​ candy, 10​% of the pieces are​ yellow, 5​% are​ red, 5​% are​ blue, 20​% are​ green, and the rest are brown.
​a) If you pick a piece at​ random, what is the probability that it is​ brown? it is yellow or​ blue? it is not​ green? it is​ striped?
​b) Assume you have an infinite supply of these candy pieces from which to draw. If you pick three pieces in a​ row, what is the probability that they are all​ brown? the third one is the first one that is​ red? none are​ yellow? at least one is​ green?

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52802)

.

8 questions in one post is TOOOOOOOOOOOOOOOOOOOOOOOOOOOO many.



Answer by math_tutor2020(3817)


I'll do part (a) to get you started.
Your teacher should have broken it up so that each part only has one question, instead of four.

---------------------------------------------------

Question: What is the probability that it is​ brown?

Solution:
10% are yellow
5% are red
5% are blue
20% are green
10+5+5+20 = 40% are either of those four colors
100-40 = 60% are brown which is the probability of selecting this color.

---------------------------------------------------

Question: What is the probability that it is​ yellow or blue?

Answer: 15%

This is because there are 10% that are yellow and 5% that are blue, so that's 10+5 = 15% total

---------------------------------------------------

Question: What is the probability it is not green?

Answer: 80%

Subtract the percentage of green (20%) from 100% to find what isn't green.

---------------------------------------------------

Question: What is the probability it is striped?

Solution:
In the information section at the top, the word "striped" is not mentioned at all. Also, there is no mention of the candies being more than one color. I'm assuming each candy is one single color only.
If this assumption is correct, then none of the candies are striped and the probability of getting a striped candy is 0%